# On the Embedding Collapse When Scaling up Recommendation Models

### v1

@@ 4.2 Is it sufficient to avoid collapse for scalability?
Evidence III: Regularization on DCNv2 to mitigate collapse.
Evidence II shows that a projection ${\bm{W}}_{i\to j}$ is learned to adjust information abundance for feature interaction. To further investigate the effect of the adjustment of ${\bm{W}}_{i\to j}$ on scalability, we introduce the following regularization with learnable parameter $\lambda_{ij}$
$\ell_{reg}=\sum_{i=1}^{N}\sum_{j=1}^{N}\left\|{\bm{W}}_{i\to j}^{\top}{\bm{W}}_{i\to j}-\lambda_{ij}{\bm{I}}\right\|_{\mathrm{F}}^{2}$
to regularize the projection matrix to be a multiplication of an unitary matrix. This way, ${\bm{W}}_{i\to j}$ will preserve all normalized singular values and maintain the information abundance after projection. We experiment with various embedding sizes and compare the changes in performance, the information abundances, and the optimization dynamics for standard and regularized models. Results are shown in Figure 5. As anticipated, regularization in DCNv2 helps learn embeddings with higher information abundance. Nevertheless, from the performance perspective, the model presents unexpected results whereby the scalability does not improve or worsen as the collapse is alleviated. We further find that such a model overfits during the learning process, with the training loss consistently decreasing and the validation AUC dropping.
(a) IA w/ 10x model size.
(b) Test AUC w.r.t. model size.
(c) Training vs. validation.
@@
Refer to caption (a) IA w/ 10x model size. Refer to caption (b) Test AUC w.r.t. model size.
@@
Evidence IV: Scaling up DCNv2 and DNN.
We now discuss DNN, which consists of a plain interaction module by concatenating all feature vectors from different fields and processing with an MLP, formulized by
${\bm{h}}=G([{\bm{e}}_{1},{\bm{e}}_{2},...,{\bm{e}}_{N}]).$
Since DNN does not conduct explicit 2-order feature interaction (Rendle et al., 2020), following our previous interaction-collapse law, it would suffer less from collapse. We compare the learned embeddings of DCNv2 and DNN and their performance with the growth of embedding size. Considering that different architectures or objectives may differ in modeling, we mainly discuss the performance trend as a fair comparison. Results are shown in Figure 6. DNN learns less-collapsed embedding matrices, reflected by higher information abundance than DCNv2. Yet, perversely, the AUC of DNN drops when increasing the embedding size, while DCNv2 sustains the performance. Such observations show that DNN falls into the issue of overfitting and lacks scalability, even though it suffers less from collapse.
@@
Regularized DCNv2 and DNN are both models with feature interaction suppressed, and they learn less-collapsed embedding matrices than DCNv2, as expected. Yet observations in evidence III&IV demonstrate that regularized DCNv2 and DNN are both non-scalable with the growth of model size and suffer from serious overfitting. We conclude the following finding:
{tcolorbox}
[colback=blue!2!white,leftrule=2.5mm,size=title] Finding 2. A less-collapsed model with feature interaction suppressed is insufficient for scalability due to overfitting concern.
@@ 5 Multi-embedding Design
In this section, we present a simple design of multi-embedding, which serves as an effective scaling design applicable to a wide range of model architecture designs. We introduce the overall architecture, present experimental results, and analyze how multi-embedding works.
5.1 Multi-Embedding for better scalability
The two-sided effect of feature interaction for scalability implies a principle for model design. That is, a scalable model should be capable of less-collapsed embeddings within the existing feature interaction framework instead of removing interaction. Based on this principle, we propose multi-embedding or ME as a simple yet efficient design to improve scalability. Specifically, we scale up the number of independent and complete embedding sets instead of the embedding size, and incorporate embedding-set-specific feature interaction layers. Similar to previous works such as group convolution (Krizhevsky et al., 2012) and multi-head attention (Vaswani et al., 2017), such design allows the model to learn different interaction patterns jointly, while a single-embedding model would be limited to the only interaction pattern that causes severe collapse. This way, the model is capable of learning diverse embedding vectors to mitigate collapse while keeping the original interaction modules. Formally, a model with $M$ sets of embeddings is defined as
@@ 5.1 Multi-Embedding for better scalability
where $m$ stands for the index of embedding set. One requirement of multi-embedding is that there should be non-linearities such as ReLU in interaction $I$; otherwise, the model is equivalent to single-embedding and hence does not capture different patterns (see Appendix B). As a solution, we add a non-linear projection after interaction for the model with linear interaction layers and reduce one MLP layer for $F$ to achieve a fair comparison. An overall architecture comparison of single-embedding and mult-embedding models with $N=2$ and $M=2$ is shown in Figure 8.
Figure 7: Architectures of single-embedding (left) and multi-embedding (right) models with $N=2$ and $M=2$.
Figure 7: Architectures of single-embedding (left) and multi-embedding (right) models with $N=2$ and $M=2$.
Figure 8: Scalability of multi-embedding on Criteo dataset.
@@ 5.2 Experiments
We conduct our experiments on two datasets for recommender systems: Criteo (Jean-Baptiste Tien, 2014) and Avazu (Steve Wang, 2014), which are large and hard benchmark datasets widely used in recommender systems. We experiment on baseline models including DNN, IPNN (Qu et al., 2016), NFwFM (Pan et al., 2018), xDeepFM (Lian et al., 2018), DCNv2 (Wang et al., 2021), FinalMLP (Mao et al., 2023) and their corresponding multi-embedding variants with 2x, 3x, 4x and 10x model size [^1] The embedding of NFwFM with 10x size on Avazu costs nearly 37.6GB memory, which exceeds our GPU memory limit. Therefore, we do not conduct 10x NFwFM on Avazu. On the other hand, the existing experiment with 4x is already sufficient for NFwFM on Avazu.. Here NFwFM is a variant of NFM He & Chua (2017) by replacing FM with FwFM. All experiments are performed with 8/1/1 training/validation/test splits, and we apply early stopping based on validation AUC. More details are shown in Appendix C.2.
Results.
We repeat each experiment 3 times and report the average test AUC with different scaling factors of the model size. Results are shown in Table 1. For the experiments with single-embedding, we observe that all the models demonstrate poor scalability. Only DCNv2 and NFwFM show slight improvements with increasing embedding sizes, with gains of 0.00036 on Criteo and 0.00090 on Avazu, respectively. For DNN, xDeepFM, and FinalMLP, which rely highly on non-explicit interaction, the performance even drops (0.00136 on Criteo and 0.00118 on Avazu) when scaled up to 10x, as discussed in Section 4.2. In contrast to single-embedding, our multi-embedding shows consistent and remarkable improvement with the growth of the embedding size, and the highest performance is always achieved with the largest 10x size. For DCNv2 and NFwFM, multi-embedding gains 0.00099 on Critio and 0.00202 on Avazu by scaling up to 10x, which is never obtained by single-embedding. Over all models and datasets, compared with baselines, the largest models averagely achieve 0.00106 improvement on the test AUC [^2] A slightly higher AUC at 0.001-level is regarded significant (Cheng et al., 2016; Guo et al., 2017; Song et al., 2019; Tian et al., 2023). Multi-embedding provides a methodology to break through the non-scalability limit of existing models. We visualize the scalability of multi-embedding on Criteo dataset in Figure 8. The standard deviation and detailed scalability comparison are shown in Appendix C.3.
@@ Results.
5.3 Analysis
Information abundance.
Multi-embedding models achieve remarkable scalability compared with single-embedding. We verify that such scalability originates from the mitigation of collapse. We compare the information abundance of single-embedding and multi-embedding DCNv2 with the 10x embedding size. As shown in Figure 9a, multi-embedding offers higher information abundance and indicates less collapsed embedding matrices.
Variations of embeddings.
Multi-embedding utilizes embedding-set-specific interactions to enrich embedding learning. We analyze the information abundance for each embedding set as shown in Figure 9b. It is observed that the embedding matrices of different sets vary in information abundance.
Different interaction patterns.
@@ 5.3 Analysis
To justify that the scalability of multi-embedding originates from different interaction patterns, we visualize $\|{\bm{W}}_{i\to j}^{(m)}\|_{\mathrm{F}}$ as the interaction pattern (Wang et al., 2021) for a multi-embedding DCNv2 model in Figure 9c. It is shown that the interaction layers learn various patterns. To further illustrate, we conduct an ablation study by restricting the divergence of $\|{\bm{W}}_{i\to j}^{(m)}\|_{\mathrm{F}}$ across all embedding sets. From results in Figure 9d, it is observed that the divergence-restricted multi-embedding model does not show similar scalability as standard multi-embedding models, indicating multi-embedding works from the diversity of interaction layers.
(a) $\mathrm{IA}({\bm{E}}_{i})$.
(b) $\mathrm{IA}({\bm{E}}_{i}^{(m)})$.
(c) $\|{\bm{W}}_{i\to j}^{(m)}\|_{\mathrm{F}}$.
(d) Restricting diversity.
Figure 9: Analysis of multi-embedding. (a): Multi-embedding learns higher information abundance. (b): Each embedding set learns diverse embeddings, relected by varying information abundance. (c): Embedding-set-specific feature interaction layers capture different interaction patterns. (d): Restricting diversity of $\|{\bm{W}}_{i\to j}^{(m)}\|_{\mathrm{F}}$ across all embedding sets leads to non-scalability.
@@ 7 Conclusion
In this paper, we highlight the non-scalability issue of existing recommendation models and identify the embedding collapse phenomenon that hinders scalability. From empirical and theoretical analysis around embedding collapse, we conclude the two-sided effect of feature interaction on scalability, i.e., feature interaction causes collapse while reducing overfitting. We propose a unified design of multi-embedding to mitigate collapse without suppressing feature interaction. Experiments on benchmark datasets demonstrate that multi-embedding consistently improves model scalability.
Appendix A Details of Toy Expermient
Appendix B Non-Linearity for Multi-Embedding

### v2

@@ 4.2 Is It Sufficient to Avoid Collapse for Scalability?
(a) IA w/ 10x size.
(b) Test AUC w.r.t. size.
(c) Training curve.
@@
(a) IA w/ 10x size.
(b) Test AUC w.r.t. size.
@@
Evidence III: Limiting the modules in interaction that leads to collapse.
Evidence I shows that a projection ${\bm{W}}_{i\to j}$ is learned to adjust information abundance for sub-embeddings and lead to collapse. We now inverstigate how surpressing such effect would result in model scalability by introducing the following regularization with learnable parameter $\lambda_{ij}$:
$\ell_{reg}=\sum_{i=1}^{N}\sum_{j=1}^{N}\left\|{\bm{W}}_{i\to j}^{\top}{\bm{W}}_{i\to j}-\lambda_{ij}{\bm{I}}\right\|_{\mathrm{F}}^{2},$
which regularizes the projection matrix to be a multiplication of an unitary matrix. This way, ${\bm{W}}_{i\to j}$ will preserve all normalized singular values and maintain the information abundance after projection. We experiment with various embedding sizes and compare the changes in performance, the information abundances, and the optimization dynamics for standard and regularized models. Results are shown in Figure 6. As anticipated, regularization in DCNv2 helps learn embeddings with higher information abundance. Nevertheless, the model presents unexpected results whereby the scalability does not improve or worsen even if the collapse is alleviated, and it is found that such a model overfits during the learning process with the training loss consistently decreasing and the validation AUC dropping.
Evidence IV: Directly avoiding explicit interaction.
We now investigate how directly suppressing the feature interaction would affect the scalability. We discuss DNN, which consists of a plain interaction module that concatenates all feature vectors from different fields and processes them with an MLP. Since DNN does not conduct explicit 2-order feature interaction (Rendle et al., 2020), it would suffer less from collapse following our previous interaction-collapse theory. We compare the learned embeddings of DCNv2 and DNN and their performance with the growth of embedding size. Considering that different architectures or objectives may differ in modeling, we mainly discuss the performance trend as a fair comparison. Results are shown in Figure 6. DNN learns less-collapsed embedding matrices, reflected by higher information abundance than DCNv2. Yet perversely, the AUC of DNN drops when increasing the embedding size. Such observations show that DNN falls into the issue of overfitting and lacks scalability, even though it suffers less from collapse.
@@
Regularized DCNv2 and DNN are both models with feature interaction suppressed, and they learn less-collapsed embedding matrices than DCNv2, as expected. Yet observations in evidence III&IV demonstrate that regularized DCNv2 and DNN both do not scale in AUC with the growth of model size and suffer from serious overfitting. We conclude the following finding:
Finding 2. A less-collapsed model with feature interaction suppressed improperly is insufficient for scalability due to overfitting concern.
@@ 5 Multi-Embedding Design
In this section, we present a simple multi-embedding design, which serves as an effective scaling mechanism applicable to a wide range of recommendation model architectures. We introduce the overall architecture, present experimental results, and analyze how multi-embedding works. We also discuss the role of data to give a comprehensive analysys for multi-embedding.
5.1 Multi-Embedding
Figure 7: Architectures of single-embedding (left) and multi-embedding (right) models.
The two-sided effect of feature interaction for scalability implies a principle for model design. That is, a scalable model should be capable of less-collapsed embeddings within the existing feature interaction framework instead of removing interaction. Based on this principle, we propose multi-embedding or ME as a simple yet efficient design to improve scalability. Specifically, we scale up the number of independent and complete embedding sets instead of the embedding size, and incorporate embedding-set-specific feature interaction modules. In line with previous works such as group convolution (Krizhevsky et al., 2012), multi-head attention (Vaswani et al., 2017), and other decoupling-based works in recommender systems (Liu et al., 2019, 2022; Weston et al., 2013), such design allows the model to learn different interaction patterns jointly and result in embedding sets with large diversity, while a single-embedding model would be limited in pattern extraction and suffer from severe collapse. With multi-embedding, the model is less influenced by the interaction-collapse theory and mitigates embedding collapse while keeping the original interaction modules. Formally, a recommendation model with $M$ sets of embeddings is defined as
@@ 5.1 Multi-Embedding
where $m$ stands for the index of embedding set. One requirement of multi-embedding is that there should be non-linearities such as ReLU within the interaction module $I$; otherwise, the model is equivalent to single-embedding and hence does not capture different patterns. [^1] See Appendix E. As a solution, we add a non-linear projection after interaction for the models with linear interaction modules and reduce one MLP layer for postprocessing module $F$ to achieve a fair comparison. An overall architecture comparison of single-embedding and mult-embedding models is shown in Figure 7.
@@ 5.2 Experiments
We conduct our experiments on two datasets for recommender systems: Criteo (Jean-Baptiste Tien, 2014) and Avazu (Steve Wang, 2014), which are large and challenging benchmark datasets widely used in recommender systems. We experiment on baseline models including DNN, IPNN (Qu et al., 2016), NFwFM (Pan et al., 2018), xDeepFM (Lian et al., 2018), DCNv2 (Wang et al., 2021), FinalMLP (Mao et al., 2023) and their corresponding multi-embedding variants with 2x, 3x, 4x and 10x model size. Here NFwFM is a variant of NFM (He & Chua, 2017) by replacing FM with FwFM. All experiments are performed with 8/1/1 training/validation/test splits, and we apply early stopping based on validation AUC. More details are shown in Appendix C.2.
@@
Results.
We repeat each experiment 3 times and report the average test AUC with different scaling factors of the model size. Results are shown in Table 1. For the experiments with single-embedding, we observe that all the models demonstrate poor scalability. Only DCNv2 and NFwFM show slight improvements with increasing embedding sizes, with gains of 0.00036 on Criteo and 0.00093 on Avazu, respectively. For DNN, xDeepFM, and FinalMLP, which rely highly on non-explicit interaction, the performance even drops (0.00136 on Criteo and 0.00118 on Avazu) when scaled up to 10x, as discussed in Section 4.2. In contrast to single-embedding, our multi-embedding shows consistent and remarkable improvement with the growth of the embedding size, and the highest performance is always achieved with the largest 10x size. For DCNv2 and NFwFM, multi-embedding gains 0.00099 on Critio and 0.00223 on Avazu by scaling up to 10x, which is never obtained by single-embedding. Over all models and datasets, compared with baselines, the largest models averagely achieve 0.00110 improvement on the test AUC [^2] A slightly higher AUC at 0.001-level is regarded significant (Cheng et al., 2016; Guo et al., 2017; Song et al., 2019; Tian et al., 2023). Multi-embedding provides a methodology to break through the non-scalability limit of existing models. We visualize the scalability of multi-embedding on Criteo dataset in Figure 8a. The standard deviation and detailed scalability comparison are shown in Appendix C.3.
(a) Multi-embedding on Criteo.
(b) $\mathrm{IA}({\bm{E}}_{i})$ on DCNv2.
Figure 8: Effectiveness of multi-embedding. (a) Model performance consistently improves with increasing size when using multi-embedding on 6 mainstream models. (b) Multi-embedding consistently enhances information abundance across all fields compared with single-embedding.
(a) $\sum_{i=1}^{N}\mathrm{IA}({\bm{E}}_{i}^{\to j})$, SE.
(b) $\sum_{i=1}^{N}\mathrm{IA}({\bm{E}}_{i}^{\to j})$, ME.
Figure 9: Information abundance of sub-embedding matrices for single-embedding and multi-embedding on DCNv2, with field indices sorted by information abundance of corresponding raw embedding matrices. Higher or warmer indicates larger. $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ are less influenced by $j$ or $\mathrm{IA}({\bm{E}}_{j})$ in multi-embedding.
Mitigation of embedding collapse.
We compare the information abundance of single-embedding and multi-embedding DCNv2 with the largest 10x embedding size to measure the mitigation of collapse, with all embedding sets for a single field concatenated together as the overall embedding in multi-embedding DCNv2. Results are shown in Figure 8b. It is observed that multi-embedding DCNv2 consistently increases the information abundance for all fields compared with single-embedding DCNv2, especially for fields with larger cardinality. These results indicate that multi-embedding is a simple yet effective method to mitigate embedding collapse for scalability gain without introducing significant computational resources or hyper-parameters.
Deployment in the online system.
After the online A/B testing in January 2023, the multi-embedding paradigm has been successfully deployed in Tencent’s Online Advertising Platform, one of the largest advertisement recommendation systems. Upgrading the click prediction model from the single-embedding paradigm to our proposed multi-embedding paradigm in WeChat Moments leads to a 3.9% GMV (Gross Merchandise Value) lift, which brings hundreds of millions of dollars in revenue lift per year. Details are introduced in Pan et al. (2024).
5.3 How Multi-Embedding works?
Less influenced by the interaction-collapse theory.
Following our previous interaction-collapse theory and corresponding analysis, embedding collapse is caused by feature interaction between different fields, reflected by the co-influence on the information abundance of sub-embeddings. We show that multi-embedding suffers less from such an effect. Recall in Section 4 that we calculate $\sum_{i=1}^{N}\mathrm{IA}({\bm{E}}_{i}^{\to j})$ to compare how $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ is influenced by the field to interact with. We here correspondingly visualize the result for multi-embedding and single-embedding DCNv2, shown in Figure 9. It can be observed that the correlation factor in multi-embedding is significantly less than that in single-embedding (0.52 vs. 0.68). Thus, the information abundance is less influenced by the field it interacts with, suffering less from the impact of the interaction-collapse theory.
(a) Diversity of ME & SE.
(b) $\|{\bm{W}}_{i\to j}^{(m)}\|_{\mathrm{F}}$.
(c) Scaling up ME variants.
(d) Diversity of ME variants.
Figure 10: Analysis of multi-embedding (ME). (a): Multi-embedding learns embedding sets with significantly larger diversity than single-embedding. (b): Embedding-set-specific feature interaction modules capture different interaction patterns. (c)(d): Using variants of ME with non-separated interaction lead to worse scalability and lower embedding set diversity.
Mitigating collapse from embedding diversity.
We further demonstrate that multi-embedding mitigates collapse by allowing the diversity of embedding sets. To illustrate, we introduce the cosine of principal angle (Miao & Ben-Israel, 1992), $\cos\left(\bm{\phi}_{i}^{m\leftrightarrow m^{\prime}}\right)$, to measure space similarity between a pair of embedding sets $m,m^{\prime}$ for any certain field $i$, calculated by the following further singular value decomposition:
$\big{(}{\bm{U}}_{i}^{(m)}\big{)}^{\top}{\bm{U}}_{i}^{(m^{\prime})}={\bm{P}}_{i}^{(m)}\mathrm{diag}\left(\cos\big{(}\bm{\phi}_{i}^{m\leftrightarrow m^{\prime}}\big{)}\right)\big{(}{\bm{P}}_{i}^{(m^{\prime})}\big{)}^{\top}.$
A low-rank $\big{(}{\bm{U}}_{i}^{(m)}\big{)}^{\top}{\bm{U}}_{i}^{(m^{\prime})}$ implies a high-rank overall embedding $[{\bm{E}}_{i}^{(m)},{\bm{E}}_{i}^{(m^{\prime})}]$. [^3] See Appendix F. We therefore discuss a gereralized measurement
$\mathrm{div}({\bm{E}}_{i}^{(m)},{\bm{E}}_{i}^{(m^{\prime})})=1-\frac{1}{K}\left\|\cos\left(\bm{\phi}_{i}^{m\leftrightarrow m^{\prime}}\right)\right\|_{1}$
to describe the diversity of embedding sets. A larger diversity implies larger information abundance for the overall embedding or better mitigation of collapse. For comparision, we split the embedding of a single-embedding DCNv2 and an ideal random-initialized matrix into embedding sets, and compare with multi-embedding DCNv2. We show the average diversity across all embedding set pairs and all fields in Figure 10a. It is shown that multi-embedding can significantly lower the embedding set similarity compared with single-embedding, mitigating embedding collapse.
Yielding diversity from separated interaction.
@@ 5.3 How Multi-Embedding works?
We further demonstrate the embedding diversity of multi-embedding models originates from embedding-set-specific feature interaction modules, which allows the embedding sets to capture diverse interaction patterns. On the one hand, we visualize $\|{\bm{W}}_{i\to j}^{(m)}\|_{\mathrm{F}}$ as the interaction pattern (Wang et al., 2021) for a multi-embedding DCNv2 model in Figure 10b. It is shown that the interaction modules learn various patterns. On the other hand, we compare multi-embedding with two variants with non-separated interaction: (a) All feature interaction modules are shared across all embedding sets, and (b) The divergence of $\|{\bm{W}}_{i\to j}^{(m)}\|_{\mathrm{F}}$ across all embedding sets are restricted by regularizations. Results are shown in Figure 10c and Figure 10d. Compared with the separated design in multi-embedding, the two variants of feature interaction design show worse scalability and embedding diversity, indicating that multi-embedding works from the separation of interaction modules.
Table 2: Averaged information abundance under various embedding sizes and data amount.
Embedding Size 5 10 15 20 25 30 35 40 40, ME
1% data 3.74 5.96 7.97 9.23 10.44 11.51 12.39 13.30 14.99
3% data 3.16 4.80 5.96 6.67 7.30 8.10 8.46 8.84 10.59
10% data 2.96 4.28 5.08 5.73 6.43 6.79 7.11 7.51 8.63
30% data 2.97 4.09 4.89 5.41 5.87 6.19 6.53 6.78 7.70
100% data 3.29 4.73 5.41 5.70 5.95 6.19 6.48 6.64 7.64
5.4 Role of Data in Embedding Collapse
Throughout our work, we mainly focus on the model scalability and conclude the intrinsic issue, embedding collapse, of recommendation models. This considerable data amount of our benchmark datasets of experiments provides credibility to the embedding collapse phenomenon within a data-amount-agnostic manner. In this section we further discuss how the embedding collapse phenomenon acts under various data amount. To illustrate, we conduct additional experiments using variously sized subsets of the Criteo dataset. We measured the average information abundance in embedding matrices across different model scales, summarized in Table 2. From the results, it is observed that the data size can indeed affect the information abundance of embedding matrices, but the information abundance does not strictly increase or even decreases along with the data size, especially for larger models. Behind this finding is that, embedding collapse is determined by two aspects: (1) the data size, which increases the information abundance, and (2) the interaction-collapse law, which decreases the information abundance. Among all the results, only for experiments with $10\%\sim 100\%$ data size and embedding size 5, and $30\%\sim 100\%$ data size and embedding size 10, 15, 20, 25, can we observe the collapse is caused by the limited data. And for most of the other cases, the abnormal decreasing trend shows that the collapse is caused by the interaction-collapse law rather than the limited data. Also throughout the results, multi-embedding consistently outperforms single-embedding under various data amount, indicating the universality of our proposed multi-embedding design.
@@ 7 Conclusion
In this paper, we highlight the non-scalability issue of existing recommendation models and identify the embedding collapse phenomenon that hinders scalability. From empirical and theoretical analysis around embedding collapse, we conclude the two-sided effect of feature interaction on scalability, i.e., feature interaction causes collapse while reducing overfitting. We propose a unified design of multi-embedding to mitigate collapse without suppressing feature interaction. Experiments on benchmark datasets demonstrate that multi-embedding consistently improves model scalability and effectively mitigates embedding collapse.
Acknowledgement
This work was supported by the National Natural Science Foundation of China (62022050 and U2342217), the BNRist Innovation Fund (BNR2024RC01010), the Tencent Innovation Fund, and the National Engineering Research Center for Big Data Software. We also gratefully acknowledge the contributions of the following: Erheng Zhong, Xueming Qiu, Cong Li, Hui Tang, Xueyu Shi and Renjie Jiang for assistance with online deployment and experiments.
Impact Statement
This paper presents work whose goal is to advance the deep learning research for scaling up recommendation models. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Criticality of Embeddings
For recommendation models, the embedding module occupies the largest number of parameters ($>92\%$ in our DCNv2 baseline for Criteo, and even larger for industrial models), and thus serves as the important and informative bottleneck part of the model. To further illustrate, we discuss the scaling up of the other modules of recommendation models, i.e., the feature interaction module $I$ and the postprocessing prediction module $F$. We experiment to increase #cross layers and #MLP layers in the DCNv2 baseline and show the results in Table 3. It is observed that increasing #cross layers or #MLP layers does not lead to performance improvement, hence it is reasonable and necessary to scale up the embedding size.
Table 3: Test AUC with enlarged feature interaction modules and portprocessing prediction modules. Higher indicates better.
DCNv2 1x 2x 4x
standard 0.81339 0.81341 0.81346
+ #cross layer 0.81325 0.81338 0.81344
+ #MLP depth 0.81337 0.81345 0.81342
Appendix B Discussion on Embeddings in Language Models
To extend our analysis to other models, we examined the pretrained T5 (Raffel et al., 2020) model, evaluating its (normalized) singular values for comparison. Results are shown in Figure 11. It is observed that T5, in contrast to DCNv2, (1) maintains higher normalized singular values and (2) exhibits a lower proportion of insignificantly small singular values, despite its larger embedding dimensions. These observations suggest that T5 is less susceptible to the embedding collapse phenomenon, possibly because text-based models are not as affected by the interaction-collapse law in field interactions that causes embedding collapse.
(a) DCNv2 vs. T5
(b) DCNv2 vs. T5 (truncated)
Figure 11: Comparison between T5 and DCNv2. Dash lines indicate the average singular values.
Appendix C Details of Experiment
C.1 Dataset Description
Table 4: Statistics of benchmark datasets for experiments.
Dataset #Instances #Fields $\sum_{i}D_{i}$ $\max\{D_{i}\}$ $\min\{D_{i}\}$
Criteo 45.8M 39 1.08M 0.19M 4
Avazu 40.4M 22 2.02M 1.61M 5
C.2 Experiment Settings
Specific multi-embedding design.
Hyperparameters.
C.3 Experimental Results
Table 5: Results on Criteo dataset. Higher indicates better.
Model Criteo
base 2x 3x 4x 10x
DNN SE 0.81228±0.00004 0.81222±0.00002 0.81207±0.00007 0.81213±0.00011 0.81142±0.00006
ME 0.81261±0.00004 0.81288±0.00015 0.81289±0.00007 0.81287±0.00005
IPNN SE 0.81272±0.00003 0.81273±0.00013 0.81272±0.00004 0.81271±0.00007 0.81262±0.00016
ME 0.81268±0.00009 0.81270±0.00002 0.81273±0.00015 0.81311±0.00008
NFwFM SE 0.81059±0.00012 0.81087±0.00008 0.81090±0.00012 0.81112±0.00011 0.81113±0.00022
ME 0.81128±0.00017 0.81153±0.00002 0.81171±0.00012 0.81210±0.00010
xDeepFM SE 0.81217±0.00003 0.81180±0.00002 0.81167±0.00008 0.81137±0.00005 0.81116±0.00009
ME 0.81236±0.00006 0.81239±0.00022 0.81255±0.00011 0.81299±0.00009
DCNv2 SE 0.81339±0.00002 0.81341±0.00007 0.81345±0.00009 0.81346±0.00011 0.81357±0.00004
ME 0.81348±0.00005 0.81361±0.00014 0.81382±0.00015 0.81385±0.00005
FinalMLP SE 0.81259±0.00009 0.81262±0.00007 0.81248±0.00008 0.81240±0.00002 0.81175±0.00020
ME 0.81290±0.00017 0.81302±0.00005 0.81303±0.00004 0.81303±0.00012
Figure 12: Visualization of scalability on Criteo dataset.
Table 6: Results on Avazu dataset. Higher indicates better.
Model Avazu
base 2x 3x 4x 10x
DNN SE 0.78744±0.00008 0.78759±0.00011 0.78752±0.00031 0.78728±0.00036 0.78648±0.00013
ME 0.78805±0.00017 0.78826±0.00013 0.78862±0.00026 0.78884±0.00005
IPNN SE 0.78732±0.00020 0.78741±0.00022 0.78738±0.00010 0.78750±0.00007 0.78745±0.00018
ME 0.78806±0.00012 0.78868±0.00023 0.78902±0.00009 0.78949±0.00028
NFwFM SE 0.78684±0.00017 0.78757±0.00020 0.78783±0.00009 0.78794±0.00022 0.78799±0.00011
ME 0.78868±0.00038 0.78901±0.00029 0.78932±0.00035 0.78974±0.00021
xDeepFM SE 0.78743±0.00009 0.78750±0.00025 0.78714±0.00030 0.78735±0.00004 0.78693±0.00050
ME 0.78848±0.00006 0.78886±0.00026 0.78894±0.00004 0.78927±0.00019
DCNv2 SE 0.78786±0.00022 0.78835±0.00023 0.78854±0.00010 0.78852±0.00003 0.78856±0.00016
ME 0.78862±0.00011 0.78882±0.00012 0.78907±0.00011 0.78942±0.00024
FinalMLP SE 0.78751±0.00026 0.78797±0.00019 0.78795±0.00017 0.78742±0.00015 0.78662±0.00025
ME 0.78821±0.00013 0.78831±0.00029 0.78836±0.00018 0.78830±0.00022
Figure 13: Visualization of scalability on Avazu dataset.
Appendix D More Baseline Methods
We also conduct experiments on AutoInt (Song et al., 2019) and compare the performance under single-embedding and multi-embedding. Due to the limited computational resource, we only scale up the model to 4x on Criteo dataset. Results are shown in Table 7. It is observed that single-embedding suffers from non-scalability while our multi-embedding consistently enhance performance along with the model size, achieving 6e-4 AUC improvement by simply scaling up.
Table 7: Results on Criteo dataset. Higher indicates better.
AutoInt base 2x 3x 4x
SE 0.81231 0.81236 0.81216 0.81193
ME 0.81264 0.81285 0.81291
Appendix E Non-Linearity for Multi-Embedding